OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = 2x^2 + 8x + 10 A. (3, -1) B. (2, -2) C. (-2, 2) D. (-1, 3) Also, if you could give an explanation that would be great! I want to be able to understand these type of problems so I don't need help in the future. Thank you!

OpenStudy (landon34):

Simplifying f(x) = 2x2 + 8x + 10 Multiply f * x fx = 2x2 + 8x + 10 Reorder the terms: fx = 10 + 8x + 2x2 Solving fx = 10 + 8x + 2x2 Solving for variable 'f'. Move all terms containing f to the left, all other terms to the right. Divide each side by 'x'. f = 10x-1 + 8 + 2x Simplifying f = 10x-1 + 8 + 2x Reorder the terms: f = 8 + 10x-1 + 2x

OpenStudy (anonymous):

haha thank you!

OpenStudy (landon34):

you need any more help, @slb4auburn ?

OpenStudy (anonymous):

actually yes, one more question

OpenStudy (landon34):

Make sure you spell every word correctly, otherwise, I cant understand It. humans are prone to doing that.

OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = -16t2 + 640t. After how many seconds does the projectile take to reach its maximum height? I think its the same concept as the other question, but I don't even know what to solve for

OpenStudy (landon34):

There is a caret or other operator missing in the equation. Example: the input 'x^2y5' is invalid because the calculator cannot determine if y should be multiplied by 5, or raised to the fifth power. To indicate that the 5 is an exponent, use the ^ symbol, otherwise use a * to indicate multiplication.

OpenStudy (landon34):

Hello?

OpenStudy (anonymous):

Thats just the question given to me, I'm not sure what to do with it

OpenStudy (anonymous):

actually, i see what you mean. it should be h(t)=-16t^2 + 640t

OpenStudy (landon34):

Simplifying h(t) = -16t2 + 640t Multiply h * t ht = -16t2 + 640t Reorder the terms: ht = 640t + -16t2 Solving ht = 640t + -16t2 Solving for variable 'h'. Move all terms containing h to the left, all other terms to the right. Divide each side by 't'. h = 640 + -16t Simplifying h = 640 + -16t

OpenStudy (landon34):

all question requested to be answered are completed, is that all you need? do you have a name?

OpenStudy (landon34):

Hello? are you there?

OpenStudy (anonymous):

that is all, thank you for your help!

OpenStudy (landon34):

you are welcome.

OpenStudy (landon34):

I must go help someone else in need now.